OpenStudy (anonymous):

1. How much energy must be absorbed by a 193-g sample of ethanol to raise its temperature from 19 to 35 C? The specific heat of ethanol is 2.4 J/g-C.

OpenStudy (chmvijay):

to increase one degree u need 2.4 J/g-C but you have temperature difference of 16 u have to increase then the energy released is =2.4*16 =X find X since its per gram of material but u have 193 gram so it will be =2.4*16*193= X *193 gram= Y ur answer

OpenStudy (anonymous):

@abb0t following chmvijays instructions I now have 38.4*193=7,411.2 I thought I was suppose to use the formula Q=m C delta T I am so confused please break it down

OpenStudy (abb0t):

yes, you should be using \(Q=mc\Delta T\)

OpenStudy (anonymous):

ok so Q=heat m=mass (in grams) C=specific heat T= the change in temperature right?

OpenStudy (abb0t):

yes.

OpenStudy (anonymous):

Q=? m=193g C=2.4 J/g*C T=19-35C whats Q? and did I fill this out right?

OpenStudy (anonymous):

@abb0t

OpenStudy (chmvijay):

ur right